A 0.79-kg ball is placed on top of a 2.5-kg book which is at the top of a table. What is the magnitude of the force exerted by the table to the two objects?
1
Expert's answer
2020-12-14T12:14:46-0500

The magnitude of the force exerted by the table to the two objects is equal to the sum of their individual weights:


"F = P_1+P_2"


By definition, the weight of a non-moving body is:


"P = mg"

where "m" is the mass of the body ("m_1 = 0.75kg" for the first and "m_2 = 2.5kg" for the second book) and "g = 9.8m\/s^2" is the gravitational acceleration. Thus, obtain:


"F = g(m_1+m_2)\\
F = 9.8\cdot (0.75+2.5) \approx 18.4N"

Answer. 18.4 N.